# Efficient Biological Data Acquisition through Inference Set Design

### v1

@@ Abstract
In drug discovery, highly automated high-throughput laboratories are used to screen a large number of compounds in search of effective drugs. These experiments are expensive, so we might hope to reduce their cost by experimenting on a subset of the compounds, and predicting the outcomes of the remaining experiments. In this work, we model this scenario as a sequential subset selection problem: we aim to select the smallest set of candidates in order to achieve some desired level of accuracy for the system as a whole. Our key observation is that, if there is heterogeneity in the difficulty of the prediction problem across the input space, selectively obtaining the labels for the hardest examples in the acquisition pool will leave only the relatively easy examples to remain in the inference set, leading to better overall system performance. We call this mechanism inference set design, and propose the use of an uncertainty-based active learning solution to prune out these challenging examples. Our algorithm includes an explicit stopping criterion that stops running the experiments when it is sufficiently confident that the system has reached the target performance. Our empirical studies on image and molecular datasets, as well as a real-world large-scale biological assay, show that deploying active learning for inference set design leads to significant reduction in experimental cost while obtaining high system performance.
@@ 1 Introduction
Automated high-throughput screening (HTS) laboratories have enabled scientists to screen large compound libraries to find effective therapeutic compounds and screen whole-genome CRISPR knockouts to understand the effects of genes on cell function (Mayr & Bojanic, 2009; Wildey et al., 2017; Blay et al., 2020; Tom et al., 2024; Fay et al., 2023). However, conducting experiments on every compound or CRISPR guide in these vast libraries remains very resource-intensive. With typical screening libraries holding on the order of $10^{5}$ to $10^{6}$ compounds (Hughes et al., 2011) and the number of possible small molecules estimated at $10^{60}$ (Bohacek et al., 1996), the disparity between our screening capabilities and all which we could explore is staggering. Reducing experimental costs without compromising the quality of the generated data would allow us to accelerate biology and pharmaceutical research and expand the set of molecules considered for testing.
@@
To avoid costs scaling with the number of perturbations, we can train a model from a subset of the target library that has been tested in the lab, and then predict experimental outcomes for the remainder of the library using the trained model (Naik et al., 2013; Reker & Schneider, 2015; Dara et al., 2022), thereby building a hybrid screen of the library. This approach entails three interrelated questions: (1) which subset of the library should we use to maximise the accuracy of the predictions?, (2) how do we select this subset without access to the experimental outcomes?, and (3) how do we ensure that we select a large enough subset to meet a target level of accuracy for the predictions?
@@
This problem is similar to an active learning problem in that we want to select examples that maximize prediction accuracy, but instead of aiming to minimize generalization error, here we only care about prediction on a particular, finite, set of experiments from a library. The fact that the library is finite introduces an important difference: the learner can influence the set of examples on which it is evaluated by strategically selecting examples. If we accept that the difficulty of predicting outcomes varies among compounds, such that some examples are inherently harder to predict, either due to their complex properties (Bengio et al., 2009), because of the partial observability of the system (Saleh et al., 2021; Krenn et al., 2020) or due to noise in the labeling function (Frénay & Verleysen, 2013; Lukasik et al., 2020), the learner can select these examples into the training set to avoid having to predict their outcomes. Conversely, if an example can be reliably predicted by the model, we can save experimental costs by not including it in the training set. We call this mechanism through which a learner can influence performance inference set design.
We propose an active learning-based solution to hybrid screening that uses the model’s confidence to guide the selection of experiments and leverages the mechanism of inference set design to improve the system’s performance on the target set. Our algorithm includes a practical stopping criterion that terminates the search for new molecules once a lower bound on a target accuracy threshold is exceeded. We show in Lemma 1 that this bound provides probabilistic guarantee on the performance of the algorithm as long as the model is weakly calibrated, such that examples for which the model is more uncertain receive a lower predicted probability than those for which it is more certain.
@@
To validate our method, we conduct series of empirical studies on image and molecular datasets, as well as a real-world case study in drug discovery. The results demonstrate inference set design significantly reduces experimental costs while improving overall system performance. Importantly, this is true even when the generalization benefits of active learning-based acquisition functions are marginal compared to random search. This has important practical implications for active learning: if a problem is a hybrid screen—in the sense that one only needs good performance on a fixed, finite set of experiments—then evaluating generalization error dramatically understates the benefits of active learning. By combining simple active learning acquisition functions with an appropriate stopping criterion, it is possible to make large scale screening far more efficient.
@@ 2 Methods
In this section, we present the problem of efficiently evaluating a large, finite set of experimental designs, such as a screening compound library. We formalize the problem as a subset selection problem, and show that it can be solved using an active learning strategy. Active learning is effective in this setting because it essentially selects its own inference set; we call this strategy, inference set design (ISD), and show that it allows to reach high levels of performance on a target set of samples. Finally, we propose a stopping criterion to monitor the model’s performance and trigger termination of the acquisition process. A pseudo-code of the algorithm is presented in Appendix A.
2.1 Problem: Minimizing Experimental Costs with Hybrid Screens
@@
We are motivated by the problem of efficient data acquisition of compound libraries for drug discovery applications. Each “experimental readout” (or label) $y$ requires us to run an experiment, which has some associated cost. We assume that we have a predefined target set $\mathcal{X}_{\text{target}}=\{x_{i}\}^{N_{\text{target}}}_{i=1}$ of experimental designs for which we want to acquire a readout. For example, in drug discovery, a common set of experimental designs, $\mathcal{X}_{\text{target}}$, would be a library of compounds, $\mathcal{D},$ to be tested on some set of cell types, $\mathcal{C}$, such that $\mathcal{X}_{\text{target}}=\mathcal{D}\times\mathcal{C}$. For each $x\in\mathcal{X}_{\text{target}}$, the associated experimental readout, $y$, would corresponds to the drug’s effect on the corresponding cell type. In this section, we assume binary labels, $y\in\{0,1\}$ for simplicity, but also explore extensions to continuous outcomes in the experiments section.
One way to reduce the acquisition costs of screening the entire target set is to train a predictive model $\hat{f}$ on some subset of the target set, $\mathcal{X}_{\text{obs}}\subseteq\mathcal{X}_{\text{target}}$, which we call the observation set, and then use this model’s predictions in place of the real labels on the remaining samples. Because experiments are typically run in multiple rounds, at each round, $t$, the target set $\mathcal{X}_{\text{target}}$ can be partitioned into two mutually exclusive subsets,
@@
  • The observation set $\mathcal{X}_{\text{obs}}^{t}$, which contains the union of all experimental designs that have already been tested in the lab and for which the readouts, $\mathcal{Y}_{\text{obs}}^{t}$, have been observed. This set grows over time as new acquisitions are made.
@@
  • The inference set $\mathcal{X}_{\text{inf}}^{t}:=\mathcal{X}_{\text{target}}\setminus\mathcal{X}_{\text{obs}}^{t}$, which consists of the remaining experimental designs in the target set that have not yet been tested. This set shrinks over time as samples are selected from it for acquisition.
@@
Once acquisition is stopped, the combination $(\mathcal{X},\mathcal{Y})_{\text{obs}}\cup(\mathcal{X},\hat{\mathcal{Y}})_{\text{inf}}$ of both the labeled pairs from the observation set and the predicted pairs from the inference set represent the output of such a system. We call this procedure a hybrid screen, providing a hybrid set of labels for the entire target set – some measured, some predicted. The system’s performance is evaluated based on its accuracy across the entire target set, which consists of perfectly accurate labels $y$ from the observation set and the inferred labels $\hat{y}$ for the inference set. Formally, the system accuracy at each round is,
$\text{ACC}_{\text{sys}}^{t}:=\frac{1}{|\mathcal{X}_{\text{target}}|}\sum_{x\in\mathcal{X}_{\text{target}}}\mathbf{1}(x\in\mathcal{X}_{\text{obs}}^{t})+\mathbf{1}(\hat{y}=y\mid x\in\mathcal{X}_{\text{inf}}^{t})$ (1)
@@
Figure 1: Hybrid screen employing active learning as an inference set design strategy. The goal is to produce a “hybrid dataset”, composed of the labels collected in the observation set and the predictions of the model made on the inference set. The system’s performance is reported as the “system accuracy”, measured as a combination of observed labels on the observation set and of predictions made on the inference set. As samples are acquired and added to the observation set, the contribution of the accuracy of the prediction model gradually decreases. The aim of the active agent is to select informative examples while pruning out the most difficult ones in order to reach high system performance without having to acquire all the samples.
@@
In this formulation, acquiring all the labels would give 100% accuracy, but at significant expense. We instead want a system that reaches an accuracy of at least $\gamma$, while incurring minimal experiment cost. This can be formulated as the following combinatorial optimization problem,
@@
$\min_{\mathcal{X}_{\text{obs}}\subseteq\mathcal{X}_{\text{target}}}|\mathcal{X}_{\text{obs}}|\quad\text{such that}\quad\text{ACC}_{\text{sys}}>\gamma$ (2)
where $|\mathcal{X}_{\text{obs}}|$ denotes the number of samples in the observation set.
@@
A natural greedy heuristic to solve this problem is to treat it as an active learning problem. Like standard active learning, we can sequentially select subsets of data from labelling and the system accuracy will benefit from any improvements in generalization that arise from better selection of training examples. However, this setting has an important difference: we only ever evaluate $\hat{f}$ on $\mathcal{X}_{\text{inf}}$, and this set is under the control of the learner because any example that is selected for labelling is removed from the inference set. As a result, by choosing acquisition functions that select harder training examples for labelling, we make the test-time task of the learner easier, because the harder examples are pruned out of the inference set. We call this approach inference set design.
@@
The data selection heuristics for inference set design are similar to standard active learning, but there are two important differences between the settings. First, if we evaluate performance under the standard active learning objective—generalization error measured on a held-out test set—then we can dramatically understate the performance improvement that result from inference set design; because of this, it is possible for active learning algorithms to dramatically outperform random search for inference set design, even when we only observe marginal improvements in generalization error. By actively selecting more challenging examples for acquisition and directing them to the observation set, the model can enhance its performance – it is allowed to actively design the inference set such that it can excel on it. This mechanism is illustrated in Figure 1. Second, unlike active learning, inference set design requires an explicit stopping criterion in order to decide when you have collected enough samples to meet the accuracy threshold.
Acquisition functions
At each timestep $t$, the model $\hat{f}^{t}$ is trained on the observation set $\mathcal{X}_{\text{obs}}^{t}$, and an acquisition function $g$ is evaluated over the entire inference set to select the next batch of samples to acquire. To design easy inference sets, we want acquisition functions that select hard examples for experimentation. For this, we evaluate both least-confidence sampling and query-by-committee. Least-confidence sampling leverages the model’s class prediction probabilities as a confidence metric, selecting samples for which the model exhibits the lowest predicted probabilities:
$\{x\}_{i=1}^{N_{b}}:=\operatorname*{arg\,max}_{x\in\mathcal{X}_{\text{inf}}}\Big{(}1-\max_{y}\,\,p(y|x)\Big{)}$ (3)
The query-by-committee approach trains an ensemble of $N_{\text{models}}$ independent predictors in parallel, selecting samples where the ensemble members exhibit the highest disagreement:
$\{x\}_{i=1}^{N_{b}}:=\operatorname*{arg\,max}_{x\in\mathcal{X}_{\text{inf}}}\,\,\mathcal{H}\Big{(}\big{\{}\operatorname*{arg\,max}_{y}\,\,p_{j}(y|x)\big{\}}_{j=1}^{N_{\text{models}}}\Big{)}$ (4)
where $\mathcal{H}$ is the entropy over the voting distribution of the ensemble members. This helps identify regions of epistemic uncertainty where model disagrement may stem from over confidence.
@@ 2.2 Empirical Stopping Criterion
The key technical challenge of inference set design is deciding when to stop running new experiments. In a production system, the labels for the inference set will remain unavailable, so we need to estimate the system performance from the collected data. To avoid unnecessary experimentation, inference set design algorithms must employ an efficient stopping criterion. We address this by maintaining a probabilistic lower bound on the system accuracy, and stopping once this bound exceeds the critical threshold, $\gamma$. A simple approach to maintaining such a bound is to leverage the feedback from each round of experimentation.
Because at every round we select examples that the model regards as challenging, the performance on each batch $\mathcal{X}_{b}^{t}$ we select at round $t$ will lower bound the performance on the inference set,
@@
Let $\hat{p}$ denote the prediction function, $\hat{f}$’s predicted probability that the outcome is 1. If the prediction function, $\hat{f}$, is approximately calibrated, such that for any $p_{1}>p_{2}$, $P[\hat{Y}=Y|\hat{p}=p_{1}]>P[\hat{Y}=Y|\hat{p}=p_{1}]$ for all $p_{1},p_{2}\in(0,1)$, then for a batch, $\mathcal{X}_{b}^{t}$, selected by least confidence, $E_{x\sim\mathcal{X}_{b}^{t}}[\mathbf{1}(Y=\hat{f}(X))]\leq E_{x\sim\mathcal{X}_{\text{inf}}^{t}}[\mathbf{1}(Y=\hat{f}(X))]$
The proof is in Appendix D. Lemma 1 shows us that we can safely use performance on $\mathcal{X}_{b}$ as a proxy for performance $\mathcal{X}_{\text{inf}}^{t}$, because the former lower-bounds the latter. Because of this, we can treat performance on $\mathcal{X}_{\text{inf}}^{t}$ as though they were IID samples from $\mathcal{X}_{\text{inf}}^{t}$, and construct a (conservative) lower bound on the accuracy on $\mathcal{X}_{\text{inf}}$. Denote the batch accuracy, $\hat{\mu}_{t}=E_{x\sim\mathcal{X}_{b}^{t}}[\mathbf{1}(Y=\hat{f}(X))]$, and its lower bound as $\alpha_{t}$. Using standard concentration inequalities (Lattimore & Szepesvári, 2020, page 135) we have that,
$P(\mu<\alpha_{t})\leq\delta\quad\text{ if }\quad\alpha_{t}=\min\left\{a\in[0,1]:KL(\hat{\mu}_{t},a)\leq\frac{\log\left(\frac{1}{\delta}\right)}{b}\right\}$ (5)
where $\mu$ denotes the accuracy on $\mathcal{X}_{\text{inf}}^{t}$. We can compute the associated lower bound on system accuracy at each round by assuming we get an accuracy of $1$ on everything in $\mathcal{X}_{\text{obs}}^{t}$ and $\alpha_{t}$ on the remaining batches. Given this, our system stops at time $\tau$ when the following criterion hold,
$\tau=\operatorname*{arg\,min}_{t\in 1,...,t}\frac{|\mathcal{X}_{\text{obs}}^{t}|+\alpha_{t}|\mathcal{X}_{\text{inf}}^{t}|}{|\mathcal{X}_{\text{target}}|}>\gamma$ (6)
which ensures achieving our desired accuracy threshold with probability at least $1-\delta$.

### v2

@@ Abstract
In drug discovery, highly automated high-throughput laboratories are used to screen a large number of compounds in search of effective drugs. These experiments are expensive, so one might hope to reduce their cost by experimenting on a subset of the compounds, and predicting the outcomes of the remaining experiments. In this work, we model this scenario as a sequential subset selection problem: we aim to select the smallest set of candidates in order to achieve some desired level of accuracy for the system as a whole. Our key observation is that, if there is heterogeneity in the difficulty of the prediction problem across the input space, selectively obtaining the labels for the hardest examples in the acquisition pool will leave only the relatively easy examples to remain in the inference set, leading to better overall system performance. We call this mechanism inference set design, and propose the use of an confidence-based active learning solution to prune out these challenging examples. Our algorithm includes an explicit stopping criterion that stops running the experiments when it is sufficiently confident that the system has reached the target performance. Our empirical studies on image and molecular datasets, as well as a real-world large-scale biological assay, show that active learning for inference set design leads to significant reduction in experimental cost while retaining high system performance.
@@ 1 Introduction
Automated high-throughput screening (HTS) laboratories have enabled scientists to screen large compound libraries to find effective therapeutic compounds and screen whole-genome CRISPR knockouts to understand the effects of genes on cell function (Mayr & Bojanic, 2009; Wildey et al., 2017; Blay et al., 2020; Tom et al., 2024; Fay et al., 2023). However, conducting experiments on every compound or gene in these vast design spaces remains very resource-intensive. With typical screening libraries holding on the order of $10^{5}$ to $10^{6}$ compounds (Hughes et al., 2011) and the number of possible small molecules estimated at $10^{60}$ (Bohacek et al., 1996), the disparity between our screening capabilities and all which we could explore is staggering. Reducing experimental costs without compromising the quality of the generated data would allow us to accelerate biology and pharmaceutical research and expand the set of molecules considered for testing.
@@
To avoid costs scaling with the number of experiments, we can train a model from a subset of the target library that has been tested in the lab, and then predict experimental outcomes for the remainder of the library using the trained model (Naik et al., 2013; Reker & Schneider, 2015; Dara et al., 2022), thereby building a hybrid screen of the library. This approach entails three interrelated questions: (1) which subset of the library should we use to maximise the accuracy of the predictions?, (2) how do we select this subset without access to the experimental outcomes?, and (3) how do we ensure that we acquire a large enough collection to meet a target level of accuracy for the predictions?
@@
This problem is similar to an active learning problem in that we want to select examples that maximize prediction accuracy, but instead of aiming to minimize generalization error, here we focus solely on prediction on a particular, finite, set of experiments from a library. The fact that the library is finite introduces an important difference: the learner can influence the set of examples on which it is evaluated by strategically selecting examples to be acquired. We assume that the difficulty of predicting outcomes varies among examples. Some experiments outcomes are inherently harder to predict either due to their complex properties (Bengio et al., 2009), because of the partial observability of the system (Saleh et al., 2021; Krenn et al., 2020) or due to noise in the labeling function (Frénay & Verleysen, 2013; Lukasik et al., 2020). Such disparity allows the learner to select these harder examples into the training set to avoid having to predict their outcomes. Conversely, if an example can be reliably predicted by the model, we can save experimental costs by not including it in the training set. We call this mechanism through which a learner can influence evaluation performance inference set design.
@@
We propose an active learning-based solution to hybrid screening that uses the model’s confidence to guide the selection of experiments and leverages the mechanism of inference set design to improve the system’s performance on the target set. Our algorithm includes a practical stopping criterion that terminates the acquisition of additional labels once a lower bound on a target accuracy threshold is exceeded, and show that this bound provides probabilistic guarantee on the performance of the algorithm as long as the model is weakly calibrated. To validate our method, we conduct a series of empirical studies on image and molecular datasets, as well as a real-world case study in drug discovery. The results demonstrate that inference set design significantly reduces experimental costs while improving overall system performance. Importantly, this is true even when the generalization benefits of active learning-based acquisition functions are marginal compared to random search. This has important practical implications for active learning: if a problem is a hybrid screen—in the sense that one only needs good performance on a fixed, finite set of experiments—then evaluating generalization error dramatically understates the benefits of active learning. By combining simple active learning acquisition functions with an appropriate stopping criterion, it is possible to make large scale screening far more efficient.
@@ 2 Methods
In this section, we present the problem of efficiently evaluating a large, finite set of experimental designs, such as a screening compound library. We formalize the problem as a subset selection problem, and show that it can be solved using an active learning strategy. Active learning is effective in this setting because it essentially selects its own inference set; we call this strategy, inference set design. We we propose a stopping criterion to monitor the model’s performance and trigger termination of the acquisition process and show that this algorithm allows to reach provably high levels of performance on a target set of samples. A pseudo-code is available in Appendix B.
2.1 Reducing Experimental Costs via Hybrid Screens and Inference Set Design
@@
We are motivated by the problem of efficient data acquisition of compound libraries for drug discovery applications. Each “experimental readout” (or label) $y$ requires us to run an experiment, which has some associated cost. We assume that we have a predefined target set $\mathcal{X}_{\text{target}}=\{x_{i}\}^{N_{\text{target}}}_{i=1}$ of experimental designs for which we want to acquire a readout, where $N_{\text{target}}$ represents the number of samples we wish to obtain labels or predictions for. For example, in drug discovery, a common set of experimental designs, $\mathcal{X}_{\text{target}}$, would be a library of compounds to be tested on some set of cell types, at different concentrations, etc. For each $x_{i}\in\mathcal{X}_{\text{target}}$, the associated experimental readout, $y_{i}$, would correspond to the drug’s effect on the corresponding cell type. In this section, we assume a classification problem but also explore extensions to continuous outcomes in Appendix E.1. One way to reduce the acquisition costs of screening the entire target set is to train a predictive model $\hat{p}(x)$ on the observation set – the subset of the target set, $\mathcal{X}_{\text{obs}}\subseteq\mathcal{X}_{\text{target}}$, for which we already have observed the labels $y$. We can then use this model’s predictions in place of the real labels on the remaining samples. Because experiments are typically run in multiple rounds, at each round $t$, the target set $\mathcal{X}_{\text{target}}$ can be partitioned into two mutually exclusive subsets:
@@
  • The observation set $\mathcal{X}_{\text{obs}}^{t}$, which contains the union of all experimental designs that have already been tested in the lab and for which the readouts, $\mathcal{Y}_{\text{obs}}^{t}$, have been observed. We denote their pairing the observation dataset $\mathcal{D}_{\text{obs}}^{t}:=(\mathcal{X},\mathcal{Y})_{\text{obs}}^{t}:=\{(x_{i},y_{i})\}_{i=1}^{N_{\text{obs}}}$. At each step $t$, a constant number of $N_{b}$ samples are acquired and added to the observation set, causing this set to grow over time.
@@
  • The inference set $\mathcal{X}_{\text{inf}}^{t}:=\mathcal{X}_{\text{target}}\setminus\mathcal{X}_{\text{obs}}^{t}$, which consists of the remaining experimental designs in the target set that have not yet been tested. This set shrinks over time as samples are selected from it for acquisition and transfered into the observation set.
@@
Once acquisition is stopped, at timestep $t=\tau$, the output of the system is a hybrid dataset, i.e. the combination $(\mathcal{X},\mathcal{Y})_{\text{obs}}^{\tau}\cup(\mathcal{X},\hat{\mathcal{Y}})^{\tau}_{\text{inf}}$ of both the labeled pairs from the observation set and the predicted pairs from the inference set. We call this procedure a hybrid screen, as it provides a hybrid set of labels for the entire target set – some measured, some predicted. The system’s performance at each step $t$ is evaluated based on its accuracy $\mu_{\text{sys}}^{t}$ across the entire target set, which consists of perfectly accurate labels $y$ from the observation set and the inferred labels $\hat{y}$ for the inference set:
$\mu_{\text{sys}}^{t}:=\frac{1}{|\mathcal{X}_{\text{target}}|}\sum_{x_{i}\in\mathcal{X}_{\text{target}}}\mathbf{1}(x_{i}\in\mathcal{X}_{\text{obs}}^{t})+\mathbf{1}(\hat{y}_{i}=y_{i}\mid x_{i}\in\mathcal{X}_{\text{inf}}^{t})$ (1)
@@
Figure 1: Hybrid screen employing active learning as an inference set design strategy. The goal is to produce a “hybrid dataset”, composed of the labels collected in the observation set and the predictions of the model made on the inference set. The system’s performance is reported as the “system accuracy”, measured as a combination of observed labels on the observation set and of predictions made on the inference set. As samples are acquired and added to the observation set, the contribution of the accuracy of the prediction model gradually decreases. The aim of the active agent is to select informative examples while pruning out the most difficult ones in order to reach very high system accuracy without having to acquire all the samples.
@@
In this formulation, acquiring all the labels would yield 100% accuracy, but at significant expense. We instead seek a system that reaches an accuracy of at least $\gamma$, while incurring minimal experimental costs. This can be formulated as the following combinatorial optimization problem,
@@
$\min_{\mathcal{X}_{\text{obs}}\subseteq\mathcal{X}_{\text{target}}}|\mathcal{X}_{\text{obs}}|\quad\text{such that}\quad\mu_{\text{sys}}^{t=\tau}>\gamma$ (2)
@@
where $|\mathcal{X}_{\text{obs}}|$ denotes the number of samples in the observation set. A natural approach is to treat this as an active learning problem. Like standard active learning, we can sequentially select subsets of data from labeling and the system accuracy will benefit from any improvements in generalization that arise from better selection of training examples. However, this setting has an important difference: we only ever evaluate $\hat{p}$ on $\mathcal{X}_{\text{inf}}$, and this set is under the control of the learner because any example that is selected for labeling is removed from the inference set. As a result, by choosing acquisition functions that select harder training examples firsts, we make the test-time task of the learner easier, as the harder examples are effectively pruned out of the inference set. We call this approach inference set design – by selective label acquisition, the agent actively designs the composition of the inference set such that it can excel on it. This mechanism is illustrated in Figure 1.
@@
The data selection heuristics for inference set design are similar to standard active learning, but there are two important differences between the settings. Firstly, if we evaluate performance under the standard active learning objective—generalization error measured on a held-out test set—we risk dramatically understating the performance improvement that result from inference set design. Instead, evaluating on the target set of interest allows to capture dramatical improvements over a random sampler even when only marginal improvements in generalization error can be observed. Secondly, unlike active learning, inference set design requires an explicit stopping criterion in order to decide when you have collected enough samples to meet the accuracy threshold, which we address in the following section.
@@ 2.2 Empirical Stopping Criterion
The key technical challenge of inference set design is deciding when to stop running new experiments. In a production system, the labels for the inference set will remain unavailable, so we need to estimate the system performance from the collected data only. To avoid unnecessary experimentation, inference set design algorithms must employ an efficient stopping criterion. We address this by maintaining a probabilistic lower bound on the system accuracy, and stopping once this bound exceeds the critical threshold, $\gamma$. A simple approach to maintaining such a bound is to leverage the feedback from each round of experimentation.
We consider a multi-class classification problem. Formally, let $\mathcal{K}:=\{1,2,\dots,K\}$ be the set of $K$ labels and $\hat{p}(x):=(\hat{p}_{1}(x),\dots,\hat{p}_{K}(x))$ be the predicted probabilities for $x$ that $y=k$ for all $k\in\mathcal{K}$. We denote $y_{i}$ the true label for a sample $x_{i}$, $\hat{y}_{i}:=\operatorname*{arg\,max}_{k}\hat{p}_{k}(x_{i})$ the predicted class by the model and $\hat{v}_{i}:=\max_{k}\hat{p}_{k}(x_{i})$ the predicted probability value for that class. We consider the use of the commonly employed least-confidence acquisition function:
$x_{\text{selected}}:=\operatorname*{arg\,max}_{x_{i}\in\mathcal{X}_{\text{inf}}\setminus\mathcal{X}_{b}}\Big{(}1-\hat{v}_{i}\Big{)}\quad\text{for}\quad i=1,\dots,N_{b}$ (3)
where $N_{b}$ is the number of samples to be acquired, i.e. the acquisition batch-size. The samples $x_{i}\in\mathcal{X}_{b}$ are selected one by one by applying the least-confidence criterion in Equation 3 iteratively to the remaining samples in $\mathcal{X}_{\text{inf}}$ that have not yet been transferred to $\mathcal{X}_{b}$. Because at every round we select examples that the model regards as challenging, if the model maintains a correct calibration ordering, the performance on a batch $\mathcal{X}_{b}^{t}$ that we select at round $t$ should lower bound the performance on the entire inference set. We express this formally in the following lemma:
@@
Assuming that $\hat{p}$’s top-predictions are weakly calibrated on the inference set, i.e. that for any two samples $x_{1}$ and $x_{2}$ in $\mathcal{D}_{\text{inf}}^{t-1}$, $\hat{v}_{1}\geq\hat{v}_{2}\implies P_{\mathcal{D}_{\text{inf}}^{t-1}}[Y=\hat{Y}|\hat{v}_{1}]\geq P_{\mathcal{D}_{\text{inf}}^{t-1}}[Y=\hat{Y}|\hat{v}_{2}]$, then the expected accuracy on a batch $\mathcal{D}_{b}^{t}:=\{(x_{i},y_{i})\}_{i=1}^{N_{b}}$ selected by least-confidence from $\mathcal{X}_{\text{inf}}^{t-1}$ is less or equal than the expected accuracy on the remaining inference set: $P_{\mathcal{D}_{b}^{t}}[Y=\hat{Y}]\leq P_{\mathcal{D}_{\text{inf}}^{t}}[Y=\hat{Y}]$.
See Appendix A.1 for more details and proof. Lemma 1 shows that although the samples in $\mathcal{X}_{b}^{t}$ are not independent and identically distributed samples from $\mathcal{X}_{\text{inf}}^{t}$, we can safely use the performance on $\mathcal{D}_{b}^{t}$ as a conservative proxy for performance on $\mathcal{D}_{\text{inf}}^{t}$. We thus leverage the measured performance on the acquired batch to estimate the performance on the remaining, unobserved inference set.
We can now use this conservative estimate to design an appropriate stopping criterion. Let $\hat{\mu}_{b}^{t}:=\frac{1}{N_{b}}\sum_{i=1}^{N_{b}}\mathbf{1}(y_{i}=\hat{y}_{i})$ denote the measured accuracy on the batch $\mathcal{D}_{b}^{t}$ at acquisition step $t$. A simple stopping criterion could consist in picking an accuracy threshold $\gamma$ for the desired system accuracy, and to stop acquiring new samples once our system accuracy estimate $\hat{\mu}_{\text{sys}}^{t}$ based on $\hat{\mu}_{b}^{t}$ surpasses $\gamma$:
$\tau=\operatorname*{arg\,min}_{t\in 1,...,T}\,\,\hat{\mu}_{\text{sys}}^{t}>\gamma\quad,\quad\hat{\mu}_{\text{sys}}^{t}=\frac{|\mathcal{X}_{\text{obs}}^{t}|+\hat{\mu}_{b}^{t}|\mathcal{X}_{\text{inf}}^{t}|}{|\mathcal{X}_{\text{target}}|}$ (4)
where $\tau$ denotes the stopping time. Optionally, we could augment this criterion with a patience hyperparameter to make it more robust to the stochasticity introduced by the finite sample size $N_{b}$. A more principled approach however is to directly account for the statistical uncertainty around $\hat{\mu}_{b}^{t}$ by using it to derive a bound $\alpha_{t}$ on the true inference accuracy $\mu_{\text{inf}}^{t}$. Using standard concentration inequalities (see detailed derivation in Appendix A.2) we have that,
$P(\mu_{\text{inf}}^{t}\leq\alpha_{t})\leq\delta\quad\text{ with }\quad\alpha_{t}=\min\left\{a\in[0,\hat{\mu}_{b}^{t}]:\text{KL}(\hat{\mu}_{b}^{t}||a)\leq\frac{\log\left(\frac{1}{\delta}\right)}{N_{b}}\right\}$ (5)
[…]
To summarize, in this section, we presented a framework for hybrid screens leveraging active learning to perform inference set design by selectively acquiring labels for hard-to-predict examples, thereby simplifying the inference set for easier prediction, and using a conservative stopping criterion to ensure that the desired level of system accuracy is achieved at stopping time. The pseudo-code for our algorithm is presented in Appendix B.